OpenStudy (anonymous):

slope 2/3 through (-9,3)

jigglypuff314 (jigglypuff314):

slope intercept form is y=mx+b m=slope so you plug in what you know (3) = (2/3)(-9) + b then solve for b

OpenStudy (anonymous):

how do you substitute and simplify

jigglypuff314 (jigglypuff314):

corrdinates of (-9,3) refer to (x,y) so x=-9 and y=3 m=slope=2/3 to simplify (3) = (2/3)(-9) + b multiply 2/3 and -9 so 3 = -6 + b then you add 6 to both sides +6 +6 9 = b so your equation is y = (2/3)x + 9

OpenStudy (anonymous):

ok thanks. might need more help.

OpenStudy (texaschic101):

or you can use point slope formula because you have the slope and a set of points.... y - y1 = m(x - x1) slope(m) = 2/3 (-9,3) x1 = -9 and y1 = 3 now lets sub our info into the formula y - y1 = m(x - x1) y - 3 = 2/3(x - (-9) y - 3 = 2/3(x + 9) y - 3 = 2/3x + 6 y = 2/3x + 6 + 3 y = 2/3x + 9 <-- your answer you can do it this way or jiggypuff's way...whichever one is easier for you :)

OpenStudy (anonymous):

the problem is find the equation of the line hat goes through (-10,-6) and is perpendicuular to y=7. I wrote y=-6. is this right.